40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog, Page 8 of 56  Not logged in ELOG logo
ID Date Author Type Category Subject
  2444   Tue Nov 12 17:44:39 2019 RaymondDailyProgressCryo vacuum chamberTank Update

Last week I moved the upper portion of the crane to the new, bolted crane support stand. Chub removed the wheeled lower section from the lab shortly thereafter. I also re-threaded the nylon lifting strap to remove slack and level the lid a bit better during lifting and moved one of the side tables next to the crane so the lid can be safely lowered after being lifted off the tank (see first photo). 

Opened the tank today to check internal dimensions. It is now closed (top bolts finger tight) but not under vacuum. The diaphragm pump was dispatched today, so will replace the dirty pump and pull vacuum again upon arrival. 

Attached a photo of the baseplate for future drill pattern reference. Note there are three anomalous holes, this is where the PEEK support poles should go. It was discovered today that these holes are 1/4-20 tapped but the PEEK pillars are dirlled/threaded for a smaller bolt.

I've attached a rough cartoon of the cold plate height relative to the optical ports and the tank wall. The outer rad shield is not shown and is slightly misaligned, but it can be easily aligned with a ~1.5 mm shim (better for thermal isolation anyways). 

Attachment 1: IMG_0258.JPG
IMG_0258.JPG
Attachment 2: IMG_0257.JPG
IMG_0257.JPG
Attachment 3: IMG_0259.JPG
IMG_0259.JPG
  2443   Tue Nov 12 03:40:39 2019 KojiLaserPD QEPD EQE vs Spot size

The QE of the (500um)^2 element has been tested with a half-power (0.51mW) instead of 0.92mW.
It is clear that the central dip depth is reduced by the lower power density.

 

Attachment 1: QE_vs_spotsize_half_power.pdf
QE_vs_spotsize_half_power.pdf
  2442   Mon Nov 11 22:19:09 2019 RaymondDailyProgressCryo vacuum chamberChanges to preserve long strap for internal chamber use

I've attached a photo of some changes to the cryocooler-tank connection design. We can save money and space by removing the 45 degree 1.33" conflat ports from the custom CH104 to 6" conflat adapter and using zero length conflat reducers at the unused 4 way cross ports, ie replace the 4-way piece blanks with holes for the vacuum line and gauge. The primary goal for these changes is to shorten the path from the cooler's heat station to the tank so that we keep the long thermal strap for use inside the tank. Also, the height is reduced from 89 cm to 59 cm.

A slightly different cupper adapter is needed to accomodate the thick strap, but no adapter will be needed anymore between the heat station and the thermal strap (same diameter round mates(new holes will need to be drilled though)).

Attachment 1: ShortAdapter.png
ShortAdapter.png
  2441   Wed Nov 6 17:05:33 2019 RaymondDailyProgressCryo vacuum chamberConflat flange tightening

Tightened all of the vacuum ports on the chamber so that the flange interfaces are all now metal-to-metal, ie full copper gasket compression. All of the ports required at least two star pattern passes before reaching this point, except for the bellows line to the turbo/backing pumps which was already at complete compression. Prior to tighening, the wide range gauge gave a pressure reading of 5.8x10-6 Torr and the ion-gauge showed 5.75x10-6 Torr. After tightening the wide range flange the reading dropped to 5.6x10-6 ; after tightening the ion-gauge flange the gauge reading dropped to 5.69x10-6

For future reference: the 4.625" flanges use 5/16" torx bolts and 1/2" nuts, and the 2.75" flanges use a 1/4" torx bolt and 7/16" nuts. 

  2440   Tue Nov 5 20:06:36 2019 RaymondUpdateCryo vacuum chamberCooler to tank connection

Attached is a drawing of the first phase (minimal vibration isolation) cryocooler attachment, where the main tank connects via the blue rimmed feedthrough. Boxed/circled components are those that will require custom fabrication:

  1. Copper adaptor from heat station to thermal strap
  2. 5.25" outer diameter 2-153 o-ring connection to conflat adapter (DN75 (4-5/8" outer diameter, 3" tube OD) pictured, but am comparing this with the DN100 (6" OD, 4" tube diameter))
  3. Conflat to ASA o-ring adapter
  4. a + b: Copper adapter from the flat strap connector to the round copper vacuum feedthrough (closeup shown in second figure)

Currently there are only two connections that require viton o-ring rather than conflat connections (cooler to piece 1, piece 3 to HV feedthrough). 

Attachment 1: CoolerCartoon.pdf
CoolerCartoon.pdf
Attachment 2: Screen_Shot_2019-11-05_at_20.25.54.png
Screen_Shot_2019-11-05_at_20.25.54.png
  2439   Fri Nov 1 12:47:18 2019 KojiLaserPD QEPD EQE vs Spot size

Clipping and saturation were investigated by the semi-analytical model. In the analysis, the waist radius of 20um at the micrometer position of 8mm is used.

1) Clipping

Firstly, the clipping loss was just geometrically calculated. Here the saturation issue was completely ignored. The elements P6, P3, and P2 have the sizes of (500um)^2, (750um)^2m, and (1000um)^2, respectively. However, these numbers could not explain the clipping loss observed at the large spot sizes. Instead, empirically the effective sizes of (350um)^2, (610um)^2, and (860um)^2 were given to match the measurement and the calculation. This is equivalent to have 70um of an insensitive band at each edge of an element (Attachment 1). These effective element sizes are used for the calculation throughout this elog entry.

2) Saturation modeling

To incorporate the saturation effect, set a threshold power density. i.e. When the power density exceeds the threshold, the power density is truncated to this threshold. (Hard saturation)

Resulting loss was estimated using numerical integration using Mathematica. When the threshold power density was set to be 0.85W/mm^2, the drop of QE was approximately matched at the waist (Attachment 2). However, this did not explain the observed much-earlier saturation at the lower density. This suggests that the saturation is not such hard.

In order to estimate the threshold power density, look at the beam size where the first saturation starts. The earlier sagging of the QE was represented by the threshold density of 0.1W/mm^2. (Attachment 3)

Attachment 1: QE_vs_spotsize_no_saturation.pdf
QE_vs_spotsize_no_saturation.pdf
Attachment 2: QE_vs_spotsize_saturation_0_85.pdf
QE_vs_spotsize_saturation_0_85.pdf
Attachment 3: QE_vs_spotsize_saturation_0_1.pdf
QE_vs_spotsize_saturation_0_1.pdf
  2438   Thu Oct 31 18:31:10 2019 KojiLaserPD QEPD EQE vs Spot size

InAsSb PD QE Test

The relationship between the spot radius and the apparent QE (EQE) was measured.

1) The spot size was checked with DataRay Beam'R2. The beam scanner was mounted on the post with a micrometer stage in the longitudinal direction. (Attachment1 upper plot)
It was confirmed that the beam is focused down to ~22um. The incident power was about 0.9mW.

2) The InAsSb detector (Sb3513A2) was mounted on the PD holder and then mounted on the stage+post. The photocurrent was amplified by a FEMTO's transimpedance amp (V/A=1e3Ohm). The dark current and the total photocurrent were measured at each measurement point with the beam aligned to the PD every time. The estimated EQEs were plotted in the lower plot of the attachment.

Note that P2, P3, and P6 elements have the size of (500um)^2, (750um)^2, and (1000um)^2, respectively.

The absolute longitudinal position of the sensor was of course slightly different from the position of the beam scanner. So the horizontal axis of the plots was arbitrary adjuted based on the symmetry.

The remarkable feature is that the QE goes down with small spot size. This is suggesting a nonlinear loss mechanism such as recombination loss when the carrier density is high.

With the present incident power, the beam size of 100um is optimal for all the element sizes. For the larger elements, a bigger beam size seems still fine.

The next step is to estimate the clipping loss and the saturation threshold with the Gaussian beam model.

Attachment 1: QE_vs_spotsize.pdf
QE_vs_spotsize.pdf
  2437   Thu Oct 24 11:58:27 2019 ChrisElectronicsCDSnew QIL AA boards

The anti-alias boards in the QIL AA chassis have been replaced with newer ones I found in the EE shop (serial numbers S1200217, S1200274, S1200275, S1200277).

The new boards (D070081-v4) have input buffers and a reasonably high input impedance (20k), unlike the old boards (D070081-v1).  However, according to the DCC revision notes, they may suffer from some excess low frequency noise caused by LT1492 opamps.  If it becomes a problem for us, we can replace those opamps.

The low input impedance of the original boards explains the anomalous ADC/DAC loopback measurement Jon made several months ago.  It should now be close to 0.5 ADC ct per DAC ct.  I have checked the DC gain for the first few channels, but have not exhaustively tested the new boards.  (Perhaps Jon has a script to automate this?)

  2436   Tue Oct 22 15:55:52 2019 KojiElectronicsGeneralBorrowed ITC510 from Cryo

From Cryo Cav setup

Borrowed ITC510 Laser Driver/TEC controller combo -> QIL

  2435   Mon Oct 21 17:07:17 2019 ChrisLab InfrastructureCryo vacuum chamberCryo chamber vacuum pressure logging

Ever since the initial pumpdown the pressure in the new cryo chamber has been stuck at ~6e-6 torr, so there's probably a small leak.

Our vacuum gauge controller has a serial communications port, which we can use to log the system pressure to aid in leak hunting. It's connected now to an unused port on fb4. A small python-based epics server queries the pressure gauges and makes the data available as two epics channels, C4:VAC-CRYO_PRES_P1 (wide range gauge) and C4:VAC-CRYO_PRES_P2 (ion gauge). These are recorded by the framebuilder. The script is stored under ~controls/services on fb4 and should start automatically on reboot.

Attachment 1: vac.png
vac.png
  2434   Fri Oct 18 14:44:22 2019 ChrisLab InfrastructureCryo vacuum chamber208V electrical service was installed in the QIL

A historical note - electricians from Facilities visited the lab several weeks ago and installed new electrical service. To do this with a minimum of disruption to the lab, they de-installed some electrical outlets along the south wall and reused the conductors. They also taped up plastic sheeting to the table enclosure to protect the squeezing and laser stabilization experiments.

  2433   Thu Oct 17 14:12:10 2019 DuoLab Infrastructure IOGEAR GWU637 failure

The spectrum analyzer SR785 uses a Ethernet-Wifi converter GWU637 from IOGEAR to connect to the WIFI in the lab. Today I am trying to download experiment data from SR785 as always but somehow it cannot find the device anymore. After struggling for a while, I restart the GWU637 device (unplug and plug the power cable) and then I can download data again.

I think it needs to be restarted after running for a couple weeks.

  2431   Fri Oct 11 17:33:47 2019 ranaDailyProgressCryo vacuum chamberClosing the vacuum chamber

we'll suffer if we use a oil based pump long term - please find and order a dry pump to back this turbo. Not only is it bad for the vac chamber, its bad for other optics in the lab,

  2430   Fri Oct 11 17:12:26 2019 Raymond, ChrisDailyProgressCryo vacuum chamberClosing the vacuum chamber

We installed the oil filter/trap on the roughing pump and began pulling a vacuum. This was delayed due to the turbo pump flashing an error message and shutting off automaticallly after failing to spin up within its preset ramp-up time (error message 1221(3)). Upon restarting the system there were no ramp up issues, likely due to the chamber already having pumped down to ~0.5 Torr at the time of restart. Fix: need to increase the turbo pump delay start time, currently at 0 (immediately spins up). After 30 minutes of pumping the pressure reached ~4e-5 Torr in the chamber. (It reached 7e-6 torr after 3 days of pumping.)

We also tested an alternate internal configuration with the unwrapped components (see attached doodle, unlabelled green disc is the cold plate). This has the advantage of thermally isolating the outer radiation shield from the cold plate, but, we found, would slightly misalign the optical input ports. 

Several ant traps were placed around the lab to combat the observed ant problem. 

Attachment 1: IMG_0204.JPG
IMG_0204.JPG
  2429   Thu Oct 10 17:28:33 2019 Raymond, ChrisDailyProgressCryo vacuum chamberRadiation shields unpacked

Today we unpacked the radiation shields and started to puzzle out how to assemble them. Attached are photos of the parts as we guessed they are intended to stack up. We didn't see how the outer shield would be supported and isolated from the cold plate, so we are contacting Rahul to clarify.

One detail not shown in these photos is the rather poor weld quality on the interior of the outer shield.

 

Attachment 1: IMG_20191010_150645.jpg
IMG_20191010_150645.jpg
Attachment 2: IMG_20191010_150700.jpg
IMG_20191010_150700.jpg
Attachment 3: IMG_20191010_150711.jpg
IMG_20191010_150711.jpg
Attachment 4: IMG_20191010_150757.jpg
IMG_20191010_150757.jpg
Attachment 5: IMG_20191010_150817.jpg
IMG_20191010_150817.jpg
Attachment 6: IMG_20191010_150826.jpg
IMG_20191010_150826.jpg
  2428   Wed Oct 9 17:40:57 2019 Raymond, Chris, DuoDailyProgressCryo vacuum chamberClosing the vacuum chamber

We closed the chamber without installing the radiation shields or cold plate in order to test the vacuum pressure of the empty system. Upon turning on the backing pump there was a fine oil mist from exhaust port, so the pump was turned off and an oil trap/filter has been purchased. 

It was estimated that 12 inch/pound of torque is required for each of the top plate bolts in order to compress the 75 Shore hardness Viton o-ring by 20% (recommended by O'Hanlon). 

 

  2427   Tue Oct 1 19:25:08 2019 JonComputingCymacsIP address changes

Someone (not me) has recently changed the IP addresses of the lab machines. I see the new assignments are the following:

10.0.1.14 ? QIL Lab fb4
10.0.1.20 ? QIL Lab qil-nfs
10.0.1.21 ? QIL Lab qil-ws1
10.0.1.22 ? QIL Lab qil-ws2
  2426   Thu Sep 26 18:05:31 2019 DuoLab Infrastructure Cryo cooler and pump test run

[Chris, Duo]
We tested the cryo cooler and the turbo pump this afternoon. We ran the cryo cooler for two hours. The equalization pressure is 275psi(the pressure before we turn it on) and operating pressure (pressure after running for two hours) is 295psi(attachment 1). The operating pressure is lower than expected; the manual indicates the pressure is expected to be 300-320psi.

We cycled the turbo pump. It appears to be functioning properly.

Attachment 1: IMG_0990.jpg
IMG_0990.jpg
  2425   Wed Sep 25 01:05:30 2019 KojiSummaryPD QEQE and dark current of InAsSb sensors

The lenses were arranged so that the spot on the PD can become smaller. A quick measurement on a (500um)^2 element showed the QE of ~80%

With the strong focusing lens of f=40mm, the beam was once expanded to a few mm. Then f=75mm lens focuses the beam to ~30um (radius). (See Attachments 1&2)

With this new beam, the QE was quickly checked. The new measurement is indicated as "Sb3513 A2P6new" in the plot. It showed the QE of ~80%.
The AOI was scanned to find any maximum, but the AOI of 0deg was the best at least with the given beam. I'm not sure yet why 500umx500um requires such small beam radius like 30um. Awesome

Attachment 1: P_20190924_233507_vHDR_On.jpg
P_20190924_233507_vHDR_On.jpg
Attachment 2: P_20190925_003614_vHDR_On.jpg
P_20190925_003614_vHDR_On.jpg
Attachment 3: InAsSb_QE.pdf
InAsSb_QE.pdf
Attachment 4: InAsSb_DarkCurrent.pdf
InAsSb_DarkCurrent.pdf
  2424   Mon Sep 23 23:48:12 2019 KojiUpdate2micronLasers2um sensor cards / focusing optics

Fiber Collimator (Thorlabs F028APC-2000+AD11F+LMR1) and MIR sensor cards (Thorlabs VRC6S Qty2) were delivered.

The sensor card is liquid crystal and seems temperature sensitive. It's slow and diffused. But at least we can now see 2um beams in a certain condition.

The fiber collimator seems working fine, but this gave me another issue. Now because the beam is small (w<500um) everywhere, I can't focus it very well. To make a focused beam, one needs a large beam, of course. Previously, the beam was not well focused. Therefore the final focused beam with f=150mm was sufficiently small like w=50um.

It looks like some kind of  telescope is necessary.

Attachment 1: IMG_8936.jpg
IMG_8936.jpg
  2423   Mon Sep 23 10:49:27 2019 KojiSummaryPD QEQE and dark current of InAsSb sensors

The QE and dark current of all the InAsSb sensors were measured. All the measurements were done in room temperature.
- The incident beam power of the 2004nm beam was 0.95mW.
- The beam was focused down to 50um gaussian radius, which was confirmed by DataRay BeamR.
- The angle of incidence was ~0deg.

- The element side (nominally Pin 2, 3, or 6) were connected to the vias boltage (negative) and the common ground was connected to the transimpedance amplifier (Shalika OP140 R=5100Ohm)
- The dark current was highly dependent on the reverse bias voltage. The QE was also bias dependent.
- Sb3512 A2 have different behavior compared to others. Alex mentioned that Sb3512 is the test batch. We can exclude this sensor from the test.
- The best QE was ~0.7 for Sb3513 A3 P2 (Pink) and Sb3513 A2 P6 (Purple). Both have the area of 500um^2. These two particular elements have low dark current of <1mA. The dark noise of this specific sensor should be measured.

Some issues of the measurements

- The transimpedance amp (TIA) has suspicious behavior. The saturation voltage was ~17V rather than <-15V. This indicates that the voltage regulators possibly have leakage of the input voltage (+/-18V) to the output line. This needs to be checked, particularly before the dark noise test.

- TIA saturation: The bias voltages could not be raised to ~1V for some PDs because of the dark noise and the saturation of the TIA. The transimpedance should be lowered by a factor of ~5.
- Because of the low bias voltages of these saturated cases, the max QEs were not reached. This also prevented from checking if there was any clipping loss. This should be checked again with the lower transimpedance.

- TBD: The angular dependence and the reflectivity of the sensor should be checked. It is difficult to carry out these tests without a sensor card.

Attachment 1: InAsSb_QE.pdf
InAsSb_QE.pdf
Attachment 2: InAsSb_DarkCurrent.pdf
InAsSb_DarkCurrent.pdf
Attachment 3: 190921_SbPD_QE.zip
  2422   Fri Sep 20 17:31:25 2019 DuoDailyProgressGeneralGeneral update

Basically two things are happening in the lab now.

1) We are assembling a shop list for the lab, buying a couple things to put the stuff together. After the purchase, relevant items will be uploaded under this post.

2) I made the TIA with OP27, trying to measure it noise. But the lab has too much environment noise. I will try measuring it after hours or in the EE shop and see if that gives a better results. Progress on that front will reply to this post later too.

  2421   Tue Sep 17 23:42:41 2019 Shalika SinghLaserPD QEMeasuring Quantum Efficiency of Extended InGaAs Photodiode

**[Internal Quantum Efficiency added]

[Koji, Shalika]

Further measurements were done after elog:2419 for Quantum Efficiency of Extended InGaAs Photodiodes(X8906). A Laser of wavelength 2um was used with an incident power of 0.80+0.02mW.  The Ophir RM9 power meter was used to check the incident power and also measure the reflectivity.

Attachment 1: The Setup. A Fibre launcher was used to project the laser along with a converging lens of the focal length of 40.0 mm which was further arranged with a subsequent converging lens of  150mm focal length. A mirror was used to reflect the laser light on the photodiode at an angle of 45o. The bias voltage was provided to pin 4 of photodiode using a Sallen Key low pass filter and the output at pin 3 of the photodiode was fed to a transimpedance amplifier (with a gain of 5.1k) which converted the photocurrent to voltage.

Attachment 2: The Quantum Efficiency is plotted with respect to different bias voltages, It was observed that the quantum efficiency increases with an increase in bias voltage. An External Quantum Efficiency of 77.4% was observed at 1V(maximum bias voltage for the photodiode). The Internal Q.E was observed to be 83.8% taking into account Reflectivity of (60.0+1) uW at an angle of 17deg. 

Attachment 3: To recreate all data

Attachment 1: IMG_8915.JPG
IMG_8915.JPG
Attachment 2: QE_X8906.pdf
QE_X8906.pdf
Attachment 3: Extended_InGaAs.zip
  2420   Tue Sep 10 18:38:18 2019 KojiNoise BudgetPD noiseDark Noise measurement of Extended InGaAs

- Previously, your TIA was pretty much dominated by the thermal noise current of the 5K transimpedance resistor (=0.129/sqrt(5000) nA/rtHz ~2pA/rtHz).
So, I believe it's impossible to measure 1pA/rtHz. Please check if you had any saturation or anything along the chain.

- Do you need SR560? If you think you are limited by the input noise of SR785 when having no SR560, you can use your whitening filter, which is supposed to be sufficient and better in terms of the output voltage range.

- Please note the serial number of the PD under the test.

- And, try to isolate your box from the optical table.

  2419   Tue Sep 10 17:17:11 2019 Shalika SinghNoise BudgetPD noiseDark Noise measurement of Extended InGaAs

**edited as per suggestions in elog:2420

The dark noise of IG22X2000T9(serial: X8906 and X8905), Extended InGaAs photodiodes was measured. A low pass sallen key filter(using OP27) with a gain of +1 and cut off frequency of 1Hz was used to provide the bias voltage to the photodiode. A transimpedance amplifier(using OPA140, refer elog:2416 for noise spectrum of TIA) with a gain of 5.1k was used to convert the output current of the photodiode to voltage. The input range was maintained at  -50 dBVpk during the measurement.

A bias voltage of 1.017 V was provided and the output voltage across the transimpedance amplifier was observed to be as follows:

X8906:    -0.030V, which implies that the dark current was -5.887uA.

X8905:   -0.097V, which implies that the dark current was -19.01uA.

Attachment 1: Setup representation

Attachment 2: Experimental Setup. It was made sure that the cables are free from any tension. Connections were made using BNC connectors. The transimpedance amplifier and sallen key filter were placed in a box and were not in direct contact with the optical bench. During measurement data was taken with a linewidth of 125mHz(was increased logarithmically for subsequent measurements, since measurement was taken in parts) with 200 averages for each set.

Attachment 3: Dark noise plot. The data was taken for X8906 for 5 different bias voltages. The input range was maintained at  -50 dBVpk during the measurement. It was observed that dark noise decreases with decrease in bias voltage.

Attachment 4: Dark noise plot. The data was taken for X8905 for 4 different bias voltages. The input range was maintained at  -46 dBVpk during the measurement. It was again observed that the dark noise decreases with a decrease in bias voltage.

***** The noise is observed very low for a 0V bias for both the photodiodes below 10kHz. It was observed that noise is high above 10kHz at all the bias voltages for both the series.

Attachment 5: Dark Current plot for both X8905 and X8906 series of photodiodes.

Attachment 6: Dark Current Density for both X8905 and X8906 series.

Although being made of the same material both the photodiodes have some difference in their dark current. It was observed that the photodiodes are very noisy at room temperature. I think they will deliver better performance at low temperatures. 

Attachment 7:  The 1/f noise was observed at 10Hz for both the series of photodiodes. 

Attachment 8:  Zip file to re-create the data.

Attachment 1: PD_test_setup.pdf
PD_test_setup.pdf
Attachment 2: setup.jpg
setup.jpg
Attachment 3: Noise_across_extended_InGaAs_X8906.pdf
Noise_across_extended_InGaAs_X8906.pdf
Attachment 4: Noise_across_extended_InGaAs_X8905.pdf
Noise_across_extended_InGaAs_X8905.pdf
Attachment 5: Dark_current_Extended_InGaAs.pdf
Dark_current_Extended_InGaAs.pdf
Attachment 6: Dark_current_density_Extended_InGaAs.pdf
Dark_current_density_Extended_InGaAs.pdf
Attachment 7: 1_fnoise_Extended_InGaAs.pdf
1_fnoise_Extended_InGaAs.pdf
Attachment 8: Extended_InGaAs.zip
  2418   Sun Sep 8 16:08:04 2019 ranaElectronicsGeneralSolder: what kind of solder to use and why?

This is a summary of some information on types of solder and their usefulness.

Summary: use the 63/37 Sn/Pb solder from Kester. It is eutectic and has a low melting point so that your opamps won't get damaged.

Eutectic:

We want our solders to be "eutectic" so that it goes from the liquid phase directly to solid with no intermediate slurry. This makes a reliable (and nice looking!) solder joint.

Lead:

The tin-lead solder is a good combo.

 

Links:

  1. Basics of solder choice from Hackaday
  2.  NASA's Tin Whiskers homepage
  3. "Tin Whiskers are Real & Complex" Maxim
  2417   Thu Sep 5 15:40:22 2019 ranaNoise BudgetPD noiseNoise Analysis of transimpedance amplifier

I'm pretty sure that the OP27 data is still not right. You should use the small binwidth and larger # of averages as we talked about earlier this week. In the elog, you should give the PSD parameters.

  2416   Thu Sep 5 11:10:08 2019 Shalika SinghNoise BudgetTIANoise Analysis of transimpedance amplifier

Noise analysis was done using SR785. SR560 was used with a flat gain of 100 to get above the noise floor of SR785. The input range was constantly maintained at -44dBVpk for all measurements. Voltage regulators LM317 and LM337 were used to power the circuit. 200 averages were taken for all the measurements. The TIA was configured with a 5.1k feedback resistor and 100pf feedback capacitor. Please refer elog:2390 for better understanding of the circuit diagram. 

** Referring to elog:2411 the 8kHz noise bump went away on its own without changing anything in the circuit. I have no clue how it happened and why it's not happening again. 

Attachment 1: Noise analysis using OP27 in transimpedance amplifier. At Frequencies below 100Hz, data was taken in 4 parts, starting from 0Hz with a span of 25Hz but with 10 number of averages(fewer averages were taken only in this case).  At high frequencies(above 100Hz) data was taken with 200 averages. A noise was observed to be 10pA/rtHz was observed at 10Hz and 3pA/rtHz above 300Hz.

Attachment 2: LT1792 was used in this case. It was seen that it is less noisy as compared to OP27. The noise was observed to be 2pA/rtHz above 20Hz.

Attachment 3: LT1012 was used for this measurement. The noise was observed to be 3pA/rtHz above 20Hz.

Attachment 4: AD820 was used for this case. The noise was observed to be 3pA/rtHz above 500Hz.

Attachment 5: OPA140 was used for the TIA during this measurement. The noise was observed to be 2pA/rtHz above 2Hz.

Attachment 6: Noise comparison between all the OpAmps used. It was seen that OP27 isn't able to deliver performance as expected because it is getting affected a lot by the noise(1/f noise). OPA140 performs better than all the others. 

Attachment 7: Zip file to re-create all data

Attachment 1: Noise_across_TIA_Op27.pdf
Noise_across_TIA_Op27.pdf
Attachment 2: Noise_across_TIA_LT1792.pdf
Noise_across_TIA_LT1792.pdf
Attachment 3: Noise_across_TIA_LT1012.pdf
Noise_across_TIA_LT1012.pdf
Attachment 4: Noise_across_TIA_AD820.pdf
Noise_across_TIA_AD820.pdf
Attachment 5: Noise_across_TIA_OPA140.pdf
Noise_across_TIA_OPA140.pdf
Attachment 6: Noise_comparison_across_TIA.pdf
Noise_comparison_across_TIA.pdf
Attachment 7: TIA_4sep.zip
  2415   Wed Sep 4 22:14:12 2019 ranaLab Infrastructure QIL lab floor plan

what about attaching a crane to the ceiling on one of the supporting beams?

 

  2414   Tue Sep 3 16:47:08 2019 DuoLab Infrastructure QIL lab floor plan

We plan to set up the big cryostat in the QIL lab. We make a plan on how to use the space in this area.

We have these items related to this experiment: the chamber, the compressor, the pump and the crane. The crane is used to lift the lid of the chamber when we open it. 

The chamber sits on the table, its diameter is about 2'4''. We put it at the corner of the table, giving us more accessible space around it.

The compressor is used to cool the system. It is connected to the chamber via the coldhead so we will need a small table to hold the coldhead at the output of the chamber.

The pump has two parts: rough pump and turbo pump. Rough pump has more noise so we put it under the table. The turbo pump is connected to the chamber and we need a stand for that too.

The current plan for the crane is that we want to screw the crane onto the floor. We do not have space for a big crane base.

 

Attachment 1: qilfloor.pdf
qilfloor.pdf
  2413   Fri Aug 30 05:11:11 2019 ChrisNoise Budget2micronLasersNoise Analysis of Voltage Regulator Circuit using SR785 Spectrum Analyzer

In case we need to seek a further reduction in the voltage regulator noise, Wenzel has kindly published their ideas for a little noise-eating circuit at the regulator output.

 

Quote:

As it was observed that normal voltage supply is noisy and not suitable for our circuit, we plan to use a voltage regulator that will help us provide a clean supply. Referring to previous elog entries the corresponding corrections were made( polarity of electrolytic capacitors, ceramic cap in parallel to electrolytic, 3V difference between input and output of respective regulators).

Attachment 1: The Circuit Diagram of Voltage regulator

The component used                   Input Voltage                              Output Voltage

a. LM7915                                    -18 V                                  -15.1 V

b. LM7815                                    18 V                                    14.86 V

c. LM317                                      18 V                                   14.96 V      

Attachment 2: Output Voltage noise of regulator circuit

The noise observed using SR785 at the output of each regulator is shown. It clearly shows that LM317 manifests less noise in comparison to LM7915 and LM7815. It will be therefore a good idea to use this to provide 15V bias in our circuit.

Attachment 3: The Scripts

Find all the scripts and data used in this measurement.

 

  2412   Thu Aug 29 15:36:49 2019 KojiNoise Budget2micronLasersNoise Bump observed at 8kHz during TIA noise analysis

You need to check the voltage noise of the regulator outputs with the opamps connected. Probably you did it. If so, it is a riddle why the 8kHz bump is not observed in the regulator outputs, but is in the opamp outputs...

Does the noise bump happen with the +/-15V supplied by 7815/7915? How about to change the capacitor values for LM317/337 to the ones recommended in the data sheet?

It is great to see the noise peaks were largely reduced by LT1792. This is what I found before although I can't explain why.

  2411   Wed Aug 28 21:25:20 2019 Shalika SinghNoise Budget2micronLasersNoise Bump observed at 8kHz during TIA noise analysis

The noise analysis for TIA was done. The circuit was in open but kept away from SR785 (to avoid any noise effect)

Attachment 1 and 2 show how the setup was placed. The wires were kept in a way that there is no tension. The wires that were used for connection from the voltage supply were twisted in order to avoid any inductance issue. The input range was kept at -44BVpk (this was maintained at all points when taking measurements with SR560) while using the SR785. SR560 was used with a flat gain of 100 in order to get above the noise of SR785 and also the AC coupling was used. LM317 and LM337 were used to provide a 15V(+/-) supply to OpAmp. The OpAmp used here is Op27. 

Attachment 3 shows the noise analysis across TIA(using Op27). It was observed that the voltage regulators help in noise reduction to a great extent at low frequencies but somehow at around 8kHz, a huge noise bump is being observed. I also checked the noise by using directly the voltage supply at the lab. It does impart high noise at low frequencies but it's clearly visible that noise bump at 8kHz isn't there. The noise bump exists only when the voltage regulators are being used with the OpAmp. I did check if the output of voltage regulators were oscillating due to some reason but they provided a constant output of 15.04V(+/-).  I did check if the OpAmp was broken but it isn't the case because the difference between the voltage at pin 2 and 3 is zero, I have two TIA on my board so I checked the noise for both of them and I observed the same results.

Attachment 4 shows the noise of TIA using LT1792. It was seen that the 8kHz noise bump is evident on even changing the OpAmp. 

I am unable to understand how is this issue coming up. I did the measurement quite a few times just to be sure It's not a one-time thing but the noise bump is dominant. 

Attachment 5: Zip

 

 

 

Attachment 1: circuit.jpg
circuit.jpg
Attachment 2: circuit_setup.jpg
circuit_setup.jpg
Attachment 3: Noise_across_TIA_op27.pdf
Noise_across_TIA_op27.pdf
Attachment 4: Noise_across_TIA_LT1792.pdf
Noise_across_TIA_LT1792.pdf
Attachment 5: TIA.zip
  2410   Mon Aug 26 10:45:43 2019 Shalika SinghNoise BudgetPD noise1/f noise analysis and dark current density

The 1/f noise and dark current density were analysed for Sb3513_A2 photodiode. 

Attachment 1: Dark current density plot

It was observed that the dark current density has a very less difference for measurements taken across 500um, 750um and 1000um. It means that the leakage current is of low magnitude.

Attachment 2: 1/f noise at 10Hz

The 1/f noise for 500um, 750um and 1000um was plotted and 1/f noise is high for 1000um as the bias is increased. and 1/f noise is high for 500um at low bias voltages.

Attachment 3: Zip file

Attachment 1: Dark_current_density.pdf
Dark_current_density.pdf
Attachment 2: 1_fnoise.pdf
1_fnoise.pdf
Attachment 3: JPL_Sb35313_A2.zip
  2409   Fri Aug 23 17:35:37 2019 KojiNoise Budget2micronLasersNoise Analysis of Circuit using SR785 Spectrum Analyser and Zero Simulation

The TF looks good. But the noise measurement is obviously limited by the SR785 noise. We need a preamp, which is only for the purpose of the measurement. It has to have the input reffered noise about a factor of a few better than the noise predicted by Zero. At high frequency, probably we will be able to use SR560. With this low noise level, probably we can just use the flat gain of 100 for the SR560 setting. This will give you the input referred noise (of the preamp) of  ~4nV/rtHz at kHz band. Note that the gain needs to be larger than 100 to have low noiseness of SR560.

Quote:

I think this is a solid measurement.

 

  2408   Fri Aug 23 14:15:58 2019 DuoNoise Budget2micronLasersNoise Analysis of Circuit using SR785 Spectrum Analyser and Zero Simulation

I think this is a solid measurement.

Quote:

The circuit has been soldered(refer entries 2399) and the noise for Sallen Key was analyzed

Attachment 1: Circuit Diagram of Sallen Key low pass filter( cut-off= 1Hz)

Attachment 2: Transfer Function of sallen key0

Attachment 3: Noise comparison between zero and SR785 measurements. The noise matches the simulated results to a great extent and also it's less noisy so can successfully be used to bias the photodiodes(1V)

Attachment 4: Zip File

 

  2407   Fri Aug 23 12:45:08 2019 Shalika SinghNoise Budget2micronLasersNoise Analysis of Circuit using SR785 Spectrum Analyser and Zero Simulation

The circuit has been soldered(refer entries 2399) and the noise for Sallen Key was analyzed

Attachment 1: Circuit Diagram of Sallen Key low pass filter( cut-off= 1Hz)

Attachment 2: Transfer Function of sallen key.  The Frequency response Measurement was done using the Swept Sine group. The input range was -50dBVpk.

Attachment 3: Noise comparison between zero and SR785 measurements. The noise matches the simulated results to a great extent and also it's less noisy so can successfully be used to bias the photodiodes(1V).

Attachment 4: Zip File

Attachment 1: sallenkeyfilter.pdf
sallenkeyfilter.pdf
Attachment 2: bodeplot_sallenkey.pdf
bodeplot_sallenkey.pdf
Attachment 3: Noise_across_SallenKey.pdf
Noise_across_SallenKey.pdf
Attachment 4: sallen.zip
  2406   Mon Aug 19 13:57:02 2019 ShalikaNoise BudgetPD noiseDark noise of Sb3513_A2

[Koji, Nathan, Duo, Shalika]

The dark noise of the Sb3513_A2 photodiode was observed using an SRS785 spectrum analyzer. Different bias voltages were provided to understand the dependency of dark noise on the bias voltage. It was observed that as the bias voltage deeps on decreasing the dark noise decrease too. A transimpedance amplifier with a gain of 5k was used to convert photocurrent into voltage. A Sallen key low pass filter was used in order to provide a low noise bias. 

** Since the gain of the TIA was 5k so the output voltage noise was divided by 5k in order to get dark current noise.

Attachment 1: Dark current noise across Sb3513_A2 500um

Attachment 2: Dark current noise across Sb3513_A2 750um

Attachment 3: Dark current noise across Sb3513_A2 1000um

Attachment 4: Dark Current across Sb3513_A2 500um, 750um, 1000um respectively

Attachment 5: zip file containing all data

[added (Nathan)] Attachment 6: Experiment setup.

Attachment 1: Noise_across_500um.pdf
Noise_across_500um.pdf
Attachment 2: Noise_across_750um.pdf
Noise_across_750um.pdf
Attachment 3: Noise_across_1000um.pdf
Noise_across_1000um.pdf
Attachment 4: Dark_current.pdf
Dark_current.pdf
Attachment 5: JPL_Sb35313_A2.zip
Attachment 6: PD_test_setup.pdf
PD_test_setup.pdf
  2405   Sun Aug 18 14:16:59 2019 DuoSummary Measuring the dark current of PD

Koji set up an experiment measuring the dark current of the photodiodes. A bias voltage is given and the current is converted to voltage via a TIA, where it is measured. Also note that in order to provide a high quality bias voltage, we LP the output of the device with a second order sallen key filter cutoff at 1Hz. 

Attachment 1: exp.pdf
exp.pdf
Attachment 2: IMG_2921.png
IMG_2921.png
  2404   Fri Aug 16 01:28:28 2019 KojiSummary2micronLasersSwitchable breakout bok

DB9 switchable breakout box is ready. We are ready to do some PD test now.

Attachment 1: P_20190815_210355.jpg
P_20190815_210355.jpg
Attachment 2: P_20190815_210419.jpg
P_20190815_210419.jpg
  2403   Thu Aug 15 15:45:30 2019 DuoLab Infrastructure Floor plan around the big cryostat

Chris and I went to the lab and made some plans about how to use the space around the optics table. Attached a drawing of it. A couple notes about the drawing:

1. Green: underneath. The rough pump is under the table. The connection to the coldhead goes on the floor.

2. Rack: electronics rack.

3. Yellow cabinet: the cabinet that has chemicals in it.

4. Turbo/Rough: pumps.

Attachment 1: autodraw_8_15_2019.png
autodraw_8_15_2019.png
  2402   Thu Aug 15 08:36:29 2019 ShalikaNoise Budget2micronLasersNoise Analysis of Voltage Regulator Circuit using SR785 Spectrum Analyzer

I found only 1uF Tanatalum Capacitors in the EE shop. Maybe a a higher capacitance of tantalum capacitor will help. 

 

Another thing, is it good to compare the input referred noise for the different types of regulators? I did a comparison between the output noise only. 

Quote:
I'm glad to see that the voltage regulators can give good noise performance. I wish I understood what the gain peaking features came from - maybe the choice of capacitors?

 

  2401   Thu Aug 15 08:03:13 2019 ranaNoise Budget2micronLasersNoise Analysis of Voltage Regulator Circuit using SR785 Spectrum Analyzer
I'm glad to see that the voltage regulators can give good noise performance. I wish I understood what the gain peaking features came from - maybe the choice of capacitors?
  2400   Thu Aug 15 02:30:29 2019 KojiSummary2micronLasersInAsSb PDs accomodated in PD housings / PD cables

InAsSb PDs were housed in the PD cages. The cages were engraved to indicate the batch (Sb3512 or 3513) and the serials (A1, A2, ...).

The PD legs does not have an indicator for the pin1. So, the tab of the PD case is directed "UP". Also the direction of the tab is marked on the cage. The tab of the short plug was also aligned to Pin1. However, the PD case is too thin and the PDs can rotate in the cases.
So the face photo was also taken so that it indicates how Pin 1 looks like from the PD face. (Attachment 4)


Also made the cable for the LaserComponents PD and the InAsSb PD. Pin n shows up as Pin n of DB9 Male connector.

Once we have the PD test is the bias circuit (with a monitor) and some patch panel kind of preparation, we can start working on the PD test.

Attachment 1: P_20190814_214159.jpg
P_20190814_214159.jpg
Attachment 2: P_20190814_214537.jpg
P_20190814_214537.jpg
Attachment 3: P_20190814_215119.jpg
P_20190814_215119.jpg
Attachment 4: P_20190814_215158.jpg
P_20190814_215158.jpg
Attachment 5: P_20190814_220035.jpg
P_20190814_220035.jpg
Attachment 6: P_20190814_220553_003.jpg
P_20190814_220553_003.jpg
Attachment 7: P_20190814_220832.jpg
P_20190814_220832.jpg
Attachment 8: P_20190814_220841.jpg
P_20190814_220841.jpg
Attachment 9: P_20190814_235137.jpg
P_20190814_235137.jpg
  2399   Wed Aug 14 19:50:37 2019 Shalika SinghNoise Budget2micronLasersNoise analysis of Sallen Key filter using SR785 and Zero simulation

The photodiode needs a 1V bias so for a clean bias we have decided to use a sallen key low pass filter with a cut off frequency at 1Hz. The quality factor of the designed sallen key filter is 0.707.

Attachment 1: The Circuit Diagram of Sallen Key filter

The gain of the circuit is 1. 

Attachment 2: The transfer function of the filter

We can see the cut off frequency at 1Hz

Attachment 3: The Input Referred Noise of filter

The input-referred voltage noise was obtained using SR785 and compared with zero simulation. It deviates a lot from the simulated results by a factor of 100.

Attachment 4: Scripts

Find all the data and scripts used for the measurements.

NOTE: Noise plot below had inconsistent units; please ignore.

 

Attachment 1: sallenkeyfilter.pdf
sallenkeyfilter.pdf
Attachment 2: Transfer_function_of_sallen_key.pdf
Transfer_function_of_sallen_key.pdf
Attachment 3: Noise_across_SallenKey.pdf
Noise_across_SallenKey.pdf
Attachment 4: sallenkey.zip
  2398   Wed Aug 14 16:07:57 2019 Nathan HollandDailyProgress2micronLasersNew AOM Driver.

After reproting phase noise issues with the VCO, here, I have changed the VCO for a high power RF amplifier - a mini circuits ZHL--5W==sma (ZHL-5W-1). It is being driven by thre Moku to provide the 80 MHz signal. Please note the following:

  • This is a high power amplifier so you should follow the correct high power amplifier turn on/off procedure.
  • The amplifier can output a maximum of 5W. However the AOM can only tolerate a maximum of 0.6 W. Ensure that an appropriately small signal is input into the amplifier, as to not break the AOM.

 

The new setup is shown in attachment 1. Attachment 2 show the phase noise of the LO arm. This is now a much more acceptable level, as compared to the previous VCO. I am outputting -19.2 dBm, and measuring 5.3 dBm on input 2. This means I am driving the AOM at 24.8 dBm, or 0.302 W.

 

With the greater drive on the AOM, as compared to previous incorrect useage of the VCO, the signal from the IFO arm is much more robust. Attachment 3 shows the phase time series of the IFO arm. You will notice the large jumps, which spoil the spectrum, not shown here. In attachment 4 I zoom into the largest of these jumps. You can see it is comprised of several linear transitions, which occur over a short time period.

 

Aidan has suggested that locking the laser is the best way to avoide these non-stationary noise sources. He suggests that these jumps could be mode hops of the laser.

Attachment 1: 20190814__Phase_Noise_Setup.pdf
20190814__Phase_Noise_Setup.pdf
Attachment 2: MachZehnder_measured_phase_noise-LO_path__20190814.pdf
MachZehnder_measured_phase_noise-LO_path__20190814.pdf
Attachment 3: MachZehnder_measured_phase-IFO_path__20190814.pdf
MachZehnder_measured_phase-IFO_path__20190814.pdf
Attachment 4: MachZehnder_measured_phase-IFO_path_zoom__20190814.pdf
MachZehnder_measured_phase-IFO_path_zoom__20190814.pdf
  2397   Tue Aug 13 19:14:04 2019 JonComputingCymacsNFS server set up

I rebuilt one of our old desktop machines to serve as NFS server for the cymac. It is running Debian 10.0 and assigned IP 10.0.1.169 (hostname qil-nfs). I installed a new 2 TB hard drive dedicated to hosting the LIGO RTS software and frame builder archive, which is shared with all other lab machines via NFS.

I have moved the new machine into the server rack and copied the contents of /opt/rtcds on the cymac into the shared location. Functionality like sitemap and the CDS tools can now be run directly from the QIL workstation (plus any other machine on which we add the NFS mount).

Attachment 1: IMG_3587.jpg
IMG_3587.jpg
  2396   Tue Aug 13 19:03:05 2019 Shalika SinghNoise Budget2micronLasersNoise Analysis of Voltage Regulator Circuit using SR785 Spectrum Analyzer

As it was observed that normal voltage supply is noisy and not suitable for our circuit, we plan to use a voltage regulator that will help us provide a clean supply. Referring to previous elog entries the corresponding corrections were made( polarity of electrolytic capacitors, ceramic cap in parallel to electrolytic, 3V difference between input and output of respective regulators).

Attachment 1: The Circuit Diagram of Voltage regulator

The component used                   Input Voltage                              Output Voltage

a. LM7915                                    -18 V                                  -15.1 V

b. LM7815                                    18 V                                    14.86 V

c. LM317                                      18 V                                   14.96 V      

Attachment 2: Output Voltage noise of regulator circuit

The noise observed using SR785 at the output of each regulator is shown. It clearly shows that LM317 manifests less noise in comparison to LM7915 and LM7815. It will be therefore a good idea to use this to provide 15V bias in our circuit.

Attachment 3: The Scripts

Find all the scripts and data used in this measurement.

Attachment 1: regulator.pdf
regulator.pdf
Attachment 2: Noise_across_Voltage_Regulator.pdf
Noise_across_Voltage_Regulator.pdf
Attachment 3: regulator.zip
  2395   Tue Aug 13 17:59:25 2019 Nathan HollandNoise Budget2micronLasersPhase Noise of Mach Zehnder with VCO.

Following up from my previous post I measured the phase noise of the Mach Zehnder setup with the AOM driven properly with its VCO. The setup is shown in attachment 1, and compressed data in attachment 2 (goto here to get the python library to decompress this data). The value of M for this data shown is +0.85 V, though I have data for other voltages - however it should affect the performance. Using the low frequency preview I was able to see that I would need to coherently subtract the phase measurements of both measurements, which restricted me to a maximum sampling frequency of 15 kHz.

The measured data is shown in attachment 3. Already you can see that the phase noise of the VCO limits the measured phase from the IFO. This also tells me that previously, when I was modulating the VCO, the AM was affecting the measured phase. When I convert this difference into frequency noise the result can be seen in attachment 4. One upside of this setup is that the signal from the IFO is much more robust, so the PLL can stay locked for longer. This is a consquency of increased drive, 24 dB more, on the AOM.

To me this demonstrates that the way forward is to replace the VCO driving the AOM with a RF amplifier, driven by a low noise (or lower noise) signal generator. The Moku is able to drive at 80 MHz. We have a Mini Circuits ZHL-5W-1 (ZHL--5W) amplifier in the laboratory. This has 46.4 dB of gain, and a maximum power output of 37 dBm. The maximum power that can be input into the AOM is 27.8 dBm (0.6 W). Thus with an appropriate setup this miniciruits amplifier should be a viable repalcement for the current VCO.

Attachment 1: 20190813__Phase_Noise_Measurement_VCO_setup.pdf
20190813__Phase_Noise_Measurement_VCO_setup.pdf
Attachment 2: MachZehnder_w_AOM_Driver_085_20190813_102449.li
Attachment 3: 20190813__MachZehnder_w_VCO-phase_noise.pdf
20190813__MachZehnder_w_VCO-phase_noise.pdf
Attachment 4: 20190813__MachZehnder_w_VCO-frequency_noise.pdf
20190813__MachZehnder_w_VCO-frequency_noise.pdf
  2394   Tue Aug 13 16:53:57 2019 Nathan HollandElectronics2micronLasersPhase Noise of Brimrose AOM Driver (VCO).

After reviewing the Brimrose AOM driver mnual, yesterday, it turns out I was previously using it incorrectly. It is a VCO with the frequency port accepting a DC voltage, between 0 and 10 V to control the frequency of the AOM - note that the mapping is not intuitive so one should refer to the manual. The modulation port is used for amplitude modulation, not frequency modulation. This port, modulation, has 50 ohm input impedance and accepts signals between DC and 10 MHz - modulating the power output. Table 1 below shows the operating parameters we should use:

Table 1: Recommended, constant, operating procedures for Brimrose AOM driver.
Port Value (V) Comments
Frequency +9.08 Gives 80 MHz modulation frequency.
Modulation +0.85 Gives 27 dBm of RF output, which is 90% of the maximum linear range of the Brimrose AOM.

 

Following this I characterised the phase noise of this VCO. Results are shown in attachment 1, for various powers. The setup is shown in figure 2, with the data for M = +0.85 V provided in attachment 3. These results show that this VCO has poor phase performance. A value of M = +0.02 V gives the same RF power as when a 0.5 Vpp signal @ 80 MHz was input into the modulation port - as I was previously doing.

This has a few implications for previous measurements:

  • The reason it was hard for the Mach Zehnder to remain locked for a substantial duration was due to the low input power to the AOM. This probably resulted in little signal in the Mach Zehnder LO arm.
  • Previously the amplitude modulation would have provided a 80 MHz signal in the Mach Zehnder LO arm

 

The script used to create the smoothed ASD can be found here.

Attachment 1: 20190813__Brimrose_AOM_Driver_phase_noise.pdf
20190813__Brimrose_AOM_Driver_phase_noise.pdf
Attachment 2: 20190813__VCO_phase_noise_setup.pdf
20190813__VCO_phase_noise_setup.pdf
Attachment 3: brimrose_AOM_driver_VCO_phase_noise_spectra.hdf5
ELOG V3.1.3-